I need help solving for y. Please explain the steps you did when doing the calculation

(3103.203+9638y)/(8904)

this is a fraction im trying to solve for y
Answer #1

Ques 3103.203 +96384 8909 Sol i cet 3103.203 + 96387 0 8904 3103:203 + 96 3&y=0 9638y = -3/03.203 y = - 3103.2013 9638 y = -